# The Cross-Lingual Cost: Retrieval Biases in RAG over Arabic-English Corpora

### v1

@@ Abstract
Our findings reveal that retrieval is a critical bottleneck in cross-lingual domain-specific scenarios, with significant performance drops occurring when the user query and supporting document languages differ. A key insight is that these failures stem primarily from the retriever’s difficulty in ranking documents across languages. Finally, we propose a simple retrieval strategy that addresses this source of failure by enforcing equal retrieval from both languages, resulting in substantial improvements in cross-lingual and overall performance. These results highlight meaningful opportunities for improving multilingual retrieval, particularly in practical, real-world RAG applications.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as the widely accepted approach for grounding large language models (LLMs) in external knowledge, with most research and development focused on high-resource languages, most notably English. However, many real-world applications, especially in corporate contexts, rely on multilingual corpora, where content spans both high- and low-resource languages. For example, internal knowledge management systems in governmental or legal domains often store content in both a high-resource language like English and the local language, while customer support systems may receive queries in the local language that require retrieving information from a corpus that mixes technical content in both languages. These scenarios introduce cross-lingual complexity, where users interact in a low-resource language while relevant information resides in a corpus containing documents in multiple languages. Prior work has shown that system performance in such cross-lingual settings tends to lag behind monolingual setups, due to challenges across both retrieval and generation Wu et al. (2024); Sharma et al. (2025); Park and Lee (2025). In this work, we focus on the English-Arabic setting – a representative and important case of high- and low-resource language interaction.
@@
Prior work has primarily focused on the generation component Liu et al. (2025); Chirkova et al. (2024), often using multilingual benchmarks derived from Wikipedia, the predominant open-domain source. However, evaluating retrieval in this context poses challenges due to several inherent characteristics: language imbalances, overlap with pretraining data, and the fact that much of Wikipedia’s knowledge is embedded in the model’s parametric memory. In contrast, our work focuses on the less explored retrieval component within a bilingual, domain-specific setting representative of real-world corporate applications.
@@
To this end, we construct benchmarks from UAE corporate datasets with parallel English-Arabic documents. Each benchmark includes a balanced set of English and Arabic queries, with answers grounded in a single language. The languages of the user query and supporting document are selected independently, enabling a systematic analysis of cross-lingual biases. Our findings show that retrieval quality is a key driver of end-to-end performance, with the cross-language setting posing the greatest challenge. Further analysis reveals that the main source of error stems not from the language mismatch between the user query and the supporting document, but from limitations in cross-lingual retrieval within a shared multilingual embedding corpus space.
@@
Finally, we propose a simple mitigation strategy tailored to the identified error source, which retrieves an equal number of passages from each language, resulting in significant improvements in cross-lingual retrieval. The effectiveness of such a basic intervention suggests that there remains substantial room for advancements in this area.
@@ 2 Related Work
Cross Lingual Information Retrieval (CLIR) is a critical capability for accessing knowledge across language boundaries, and has gained renewed attention with the rise of cross-lingual retrieval-augmented generation (RAG) systems. These systems typically operate in two phases, retrieval and answer generation. CLIR has historically been done via translation (see Galuščáková et al. (2021) and references within). With the rise of dense retrieval, most leading techniques avoid direct translation and instead embed queries and documents of different languages into the same space Chen et al. (2024); Louis et al. (2025); Wang et al. (2024); Asai et al. (2021b). The improved performance over retrieval tasks was also verified to occur in RAG for question answering by Chirkova et al. (2024) that show an advantage to these direct methods over translation coupled with monolingual retrieval. The different retrieval techniques vary in their training method and data collections, yet all follow the same pattern of embedding the query and the document. They fall into the broader area of Cross Lingual Allignment, where the objective is to align representations of different languages Hämmerl et al. (2024). This broader area and the specifics of the different models are outside the scope of this paper.
@@ 3 Evaluation Pipeline
Our RAG pipeline consists of the standard components: retrieval, re-ranking, and answer generation. Retrieval is performed using dense vector search over a bilingual corpus. We experiment with the multilingual embedding models BAAI BGE-M3111https://huggingface.co/BAAI/bge-m3 (referred to as BGE-M3 from now on) and Multilingual-E5-Large222intfloat/multilingual-e5-large (referred to as M-E5), both of dimension $10241024$, along with the BGE-v2-M3333https://huggingface.co/BAAI/bge-reranker-v2-m3 re-ranker. These models were chosen because they are considered among the top-performing open-source retrievers and re-ranker, and have been widely adopted in previous works. For answer generation, we use Qwen-2.5-14B-Instruct [^4] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct, a large language model with strong multilingual capabilites. During inference, the 20 most relevant passages are retrieved for a given question, then re-ranked based on their relevance and utility for answer generation. The top-5 ranked passages are used to augment the prompt provided to the LLM for answer generation (using Prompt A.3).
@@ 3.1 Metrics
The end-to-end performance and the generation component are evaluated using an answer quality metric, which we refer to as accuracy, based on a semantic equivalence to ground-truth answers provided by our benchmarks (see next section). Specifically, we adopt an LLM-as-a-judge approach to assess correctness, using Claude 3.5 Sonnet to determine whether a generated answer matches the ground-truth reference (see Prompt A.3), following recent work by Zheng et al. (2023). Although LLM-based judgments have faced critique, particularly for relevance assessment Soboroff (2024), prior studies have shown a high correlation with human evaluations in QA contexts. Moreover, the common alternative of strict lexical match is even less reliable in a multilingual setting, as discussed for example in Qi et al. (2025), making a semantic measure more appropriate.
@@
For evaluating the retrieval component, we measure whether the ground-truth answer can be inferred from each retrieved passage. We obtain these relevance judgments using Claude 3.5 Sonnet with Prompt A.3. Based on these relevance labels, we report Hits@20, indicating whether a relevant passage appears among the top $2020$ retrieved results. For reranking, we apply the same procedure and report Hits@5 to measure whether relevant passages appear among the top results of the reranked list. Measuring the presence of relevant passages among the top results is particularly important in a RAG setting, as it reflects whether downstream components have access to the required evidence. The validity of these metrics is supported by their correlation with downstream accuracy, as detailed in Appendix A.1.2.
@@ 3.2 Our Benchmarks
Having each document available in both languages is essential for our experimental design. In order to build a corpus for each of these two use cases, we assign a document language to each document uniformly at random during corpus construction, ensuring that every document appears in exactly one language within the corpus. The resulting Legal corpus includes roughly 1.5M words, while the Travel corpus contains around 150K words. After building and indexing this bilingual corpus, we proceed to create the benchmark. We used DataMorgana Filice et al. (2025), a synthetic question–answer generation tool, to create query–answer pairs per document, ensuring that each question could be answered using that document alone. The language of each query–answer pair (the user language) is also selected uniformly at random and independently of the document language, resulting in a benchmark that supports systematic evaluation across all language combinations, and allows to identify the source of bias. The final benchmarks include around 1.6K question–answer pairs for Legal and 2K for Travel. Details of the DataMorgana configuration we used to generate our benchmarks are provided in Appendix A.2777The benchmarks will be made publicly available upon publication of this paper..
@@ 4 Experiments
We present four experiments, each structured with a description, results, and key conclusions. The first experiment demonstrates that retrieval is a major bottleneck in our bilingual setting. The second identifies the root cause of the retrieval issues as the cross-lingual mismatch between the user language and the document language. The third experiment attributes the performance drop to the presence of a bilingual corpus. Finally, the fourth experiment proposes and evaluates a mitigation strategy to address this issue.
@@ 4.1 Retrieval is a Critical Bottleneck
Looking more closely at the individual components, the reranker performs the best of the three. For both benchmarks with the BGE-M3 embedder, the probability of retrieval failure is comparable to that of generation. With the M-E5 embedder, the retrieval gap is even larger than the generation gap, showing a 12% difference on the Legal benchmark and 9% on Travel. Moreover, for each benchmark, reranking and generation performance are stable across embedders. However, changing retrievers has a significant effect on end-to-end accuracy. These results, taken together, highlight that the retriever is a critical bottleneck and motivate us to focus our efforts on it.
@@ 4.2 Cross-Lingual Combinations are the Most Challenging
Benchmark User Lang. Doc Lang. Retrieval Hit@20 End-to-End Accuracy
Legal Arabic Arabic 92±3% 67±5%
Arabic English 89±3% 68±4%
English Arabic 56±5% 30±5%
English English 86±3% 68±4%
[…]
Cross-lang. 72±3% 49±3%
@@
Arabic English 91±2% 78±4%
English Arabic 79±4% 72±4%
English English 93±2% 84±3%
[…]
Cross-lang. 85±2% 75±3%
@@
Benchmark User Lang. Doc Lang. Retrieval Hit@20 End-to-End Accuracy
Legal Arabic Arabic 88±4% 67±5%
Arabic English 51±4% 38±4%
@@
English English 88±3% 70±4%
[…]
Cross-lang. 46±3% 30±3%
@@
Arabic English 55±4% 38±4%
English Arabic 64±4% 62±4%
English English 95±2% 86±3%
[…]
Cross-lang. 59±3% 50±3%
@@
Table 2: Performance across language combinations. Results are presented for each embedder, benchmark and for each of the four possible user–document language combinations. In addition, we report same-language and cross-language scores, defined as the mean scores over combinations where the user and document languages match or differ, respectively.
@@
The tables reveal that cross-lingual scenarios, where the user query and the supporting document are in different languages, consistently underperform compared to same-language settings in both retrieval and end-to-end performance. For the BGE-M3 embedder, a significant decline in retrieval performance is observed only when the user language is English and the document language is Arabic, with drops of 33% in the Legal benchmark and 14% in Travel compared to the same-language configuration. A similar pattern appears in the final accuracy, with decreases of 38% and 12%, respectively. Notably, the reverse cross-lingual setting does not exhibit any statistically significant degradation for BGE-M3.
In contrast, the M-E5 embedding exhibits an even larger performance drop across both cross-lingual settings. Specifically, retrieval Hit@20 decreases by 42% on the Legal benchmark and by 33% on Travel, compared to their same-language counterparts. These retrieval declines also propagate to the end-to-end accuracy, resulting in drops of 38% for Legal and 36% for Travel.
@@ 4.3 The Source of the Cross-Lingual Failure
Notice that in our current setup, referred to from now on as the direct setting, multilinguality arises from both the corpus and the queries. This creates two main challenges. First, the retriever must handle cross-lingual queries, ranking documents in language X given a query in language Y. Second, it must rank documents across languages without introducing bias toward either a high-resource language or the user language.
To determine which of these challenges is primarily responsible for the observed failures, we conducted the following experiment. We modified the retriever from the direct setting to search only within the correct language. Specifically, for a document language X, the language-oracle retriever returns the top results exclusively in language X, completely excluding language Y. Hence, while the user query can still be in a different language, all retrieved passages, and therefore all subsequent stages of re-ranking and generation, operate exclusively on monolingual documents. This setup allows us to isolate the challenges of multilingual query embeddings while avoiding issues related to multilingual document embeddings. We stress that the oracle is used only for analysis purposes, since in practice we do not have access to the document language ahead of time. The first two bars in each subfigure of Figure 1 present the Hit@20 performance of the direct and language-oracle retrievers, broken down by user-document language combinations, as well as overall.
@@
We observe two clear phenomena. First, the language-oracle retriever achieves nearly identical performance across all user and document language pairs, suggesting there are essentially no failures related to the query embeddings. In contrast, the gap between the direct and language-oracle retrievers can be substantial in many cross-lingual cases. This indicates that the main source of failure lies in the multilingual document embeddings, specifically the retriever’s ability to rank documents across languages.
The results suggests that while semantic similarity is well captured within a single language, the embeddings struggle in cross-lingual settings. For instance, BGE-M3 appears to favor English passages when the user query is in English, while M-E5 may exhibit a tendency to prefer passages in the same language as the user query.

### v2

@@ Abstract
Our findings reveal that retrieval is a critical bottleneck in cross-lingual domain-specific scenarios, with substantial performance drops occurring when the user query and supporting document languages differ. A key insight is that these failures stem primarily from the retriever’s difficulty in ranking documents across languages. Finally, we propose two simple retrieval strategies that address this source of failure by enforcing equal retrieval from both languages or by translating the query, resulting in substantial improvements in cross-lingual and overall performance. These results highlight meaningful opportunities for improving multilingual retrieval, particularly in practical, real-world RAG applications.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as the widely accepted approach for grounding large language models (LLMs) in external knowledge, with most research and development focused on high-resource languages, most notably English. However, many real-world applications, especially in corporate contexts, rely on multilingual corpora, where content spans both high- and low-resource languages. For example, internal knowledge management systems in governmental or legal domains often store content in both a high-resource language like English and the local language, while customer support systems may receive queries in the local language that require retrieving information from a corpus that mixes technical content in both languages. These scenarios introduce cross-lingual complexity, where users interact in a low-resource language while relevant information resides in a corpus containing documents in multiple languages. Prior work has shown that system performance in such cross-lingual settings tends to lag behind monolingual setups, due to challenges across both retrieval and generation Wu et al. (2024); Sharma et al. (2025); Park and Lee (2025). In this work, we focus on the bi-lingual English-Arabic setting – a representative and important case of high- and low-resource language interaction.
@@
Prior work has primarily focused on the generation component Liu et al. (2025); Chirkova et al. (2024), often using multilingual benchmarks derived from Wikipedia, the predominant open-domain source. However, evaluating retrieval in this context poses challenges due to several inherent characteristics: language imbalances, overlap with pretraining data, and the fact that much of Wikipedia’s knowledge is embedded in the model’s parametric memory. In contrast, our work focuses on the less explored retrieval component within a bilingual, domain-specific setting representative of real-world corporate applications. In this context, we study retrieval bias, namely the tendency of multilingual retrievers to favor one language over another, thereby overlooking relevant documents in the less-preferred language. In particular, we examine the cross-lingual setting, in which a query in one language may be answered by a document written in another.
@@
We construct benchmarks from UAE corporate datasets with parallel English-Arabic documents. Each benchmark includes a balanced set of English and Arabic queries, with answers grounded in a single language. The languages of the user query and supporting document are selected independently, enabling a systematic analysis of cross-lingual biases. Our analysis of these two benchmarks highlights that retrieval presents a significant bottleneck within the RAG pipeline. Moreover, the primary source of retrieval error arises in cross-lingual settings, namely when the query and the ground truth document are in different languages.
@@
Finally, we propose two simple mitigation strategies tailored to the identified error source. The first strategy selects an equal number of passages from each language-specific subset, while the second searches the joint dataset twice, once with the original query and once with its translation. Both strategies result in substantial improvements in cross-lingual retrieval. The effectiveness of such basic interventions suggests that there remains considerable room for advancements in this area.
@@ 2 Related Work
Cross Lingual Information Retrieval (CLIR) is a critical capability for accessing knowledge across language boundaries, and has gained renewed attention with the rise of cross-lingual retrieval-augmented generation (RAG) systems. These systems typically operate in two phases, retrieval and answer generation. CLIR has historically been done via translation (see Galuščáková et al. (2021) and references within). With the rise of dense retrieval, most leading techniques avoid direct translation and instead embed queries and documents of different languages into the same space Chen et al. (2024b); Louis et al. (2025); Wang et al. (2024); Asai et al. (2021b). The improved performance over retrieval tasks was also verified to occur in RAG for question answering by Chirkova et al. (2024) that show an advantage to these direct methods over translation coupled with monolingual retrieval. The different retrieval techniques vary in their training method and data collections, yet all follow the same pattern of embedding the query and the document. They fall into the broader area of Cross Lingual Alignment, where the objective is to align representations of different languages Hämmerl et al. (2024). This broader area and the specifics of the different models are outside the scope of this paper.
@@ 3 Evaluation Pipeline
Our RAG pipeline consists of the standard components: retrieval, re-ranking, and answer generation. Retrieval is performed using dense vector search over a bilingual corpus [^1] We split documents into passages, using LlamaIndex’s SentenceSplitter into passage of up to 100 tokens with no overlap. To preserve context, each passage retained the original document title, which corresponds to the law in the Legal benchmark and to the country in the Travel benchmark.. We experiment with the multilingual embedding models BAAI BGE-M3222https://huggingface.co/BAAI/bge-m3 (referred to as BGE-M3 from now on) and Multilingual-E5-Large [^3] https://huggingface.co/intfloat/multilingual-e5-large (referred to as M-E5), both of dimension $1024$, along with the BGE-v2-M3444https://huggingface.co/BAAI/bge-reranker-v2-m3 re-ranker. These models were chosen for their recency, popularity, and status as top-performing open-source retrievers and re-rankers Li et al. (2023); Chen et al. (2024a); Wang et al. (2024); Enevoldsen et al. (2025).
For answer generation, we use Qwen-2.5-14B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct, a generative language model with strong multilingual capabilities, and part of the Qwen-2 family (Yang et al., 2024). During inference, the 20 most relevant passages are retrieved for a given question, then re-ranked based on their relevance and utility for answer generation. The top-5 ranked passages are used to augment the prompt provided to the LLM for answer generation (using Prompt A.3).
@@ 3.1 Metrics
The end-to-end performance and the generation component are evaluated using an answer quality metric, which we refer to as accuracy, based on a semantic equivalence to ground-truth answers provided by our benchmarks (see Section 3.2). Specifically, we adopt an LLM-as-a-judge approach to assess correctness, using Claude 3.5 Sonnet to determine whether a generated answer matches the ground-truth reference (see Prompt A.3), following recent work by Zheng et al. (2023). Although LLM-based judgments have faced critique, particularly for relevance assessment Soboroff (2024), prior studies have shown a high correlation with human evaluations in QA contexts. Moreover, the common alternative of strict lexical match is even less reliable in a multilingual setting, as discussed for example in Qi et al. (2025), making a semantic measure more appropriate.
@@
For evaluating the retrieval component, we measure whether the ground-truth answer can be inferred from each retrieved passage. We obtain these relevance judgments using Claude 3.5 Sonnet with Prompt A.3. Based on these relevance labels, we report Hits@20, indicating whether a relevant passage appears among the top $20$ retrieved results. For reranking, we apply the same procedure and report Hits@5 to measure whether relevant passages appear among the top results of the reranked list. Measuring the presence of relevant passages among the top results is particularly important in a RAG setting, as it reflects whether downstream components have access to the required evidence. The validity of these metrics is supported by their correlation with downstream accuracy, as detailed in Appendix A.1.2. Finally, to demonstrate that the Hit@20 results are consistent with other common metrics, Appendix A.4 also reports the NDCG and MRR corresponding to the results presented in this paper.
@@ 3.2 Our Benchmarks
Having each document available in both languages is essential for our experimental design. In order to build a corpus for each of these two use cases, we assign a document language to each document uniformly at random during corpus construction, ensuring that every document appears in exactly one language within the corpus. The resulting Legal corpus includes roughly 1.5M words, while the Travel corpus contains around 150K words. After building and indexing this bilingual corpus, we proceeded to create the benchmark. We used DataMorgana Filice et al. (2025), a synthetic question–answer generation tool, to create query–answer pairs per document, ensuring that each question could be answered using that document alone. The language of each query–answer pair (the user language) is also selected uniformly at random and independently of the document language, resulting in a benchmark that supports systematic evaluation across all language combinations, and allows to identify the source of bias. The final benchmarks include around 1.3K question–answer pairs for Legal and 2K for Travel. Details of the DataMorgana configuration we used to generate our benchmarks, along with basic statistics, are provided in Appendix A.2888Our benchmarks and corpora are available at: https://github.com/chenamiraz/cross-lingual-cost. In the Legal index, the law id serves as the document id, while in the Travel index, the country name is used as the title. .
@@ 4 Experiments
We present four experiments, each structured with a description, results, and key conclusions. The first experiment demonstrates that retrieval is a major bottleneck in our bilingual setting. The second reveals performance gaps between same-language and cross-lingual cases, with substantially worse results when the user and document languages differ. The third attributes this performance drop to the retriever’s need to rank documents in both languages against the query simultaneously. Finally, the fourth proposes and evaluates mitigation strategies to address this issue.
@@ 4.1 Retrieval is a Critical Bottleneck
Looking more closely at the individual components, the reranker performs the best of the three. For both benchmarks with the BGE-M3 embedder, the probability of retrieval failure is comparable to that of generation. With the M-E5 embedder, the retrieval gap is even larger than the generation gap, showing a 12% difference on the Legal benchmark and 9% on Travel. Moreover, for each benchmark, reranking and generation performance are stable across embedders. However, changing retrievers has a substantial effect on end-to-end accuracy. These results, taken together, highlight that the retriever is a critical bottleneck and motivate us to focus our efforts on it.
@@ 4.2 Cross-Lingual Combinations are the Most Challenging
Benchmark
User
Lang.
Doc
Lang.
Retrieval
Hit@20
End-to-End
Accuracy
Legal Arabic Arabic 92±3% 68±5%
Arabic English 90±3% 67±5%
English Arabic 56±5% 31±5%
English English 86±4% 68±5%
Same-lang. 89±2% 68±4%
Cross-lang. 73±3% 49±4%
@@
Arabic English 91±3% 78±4%
English Arabic 80±4% 70±4%
English English 94±2% 84±3%
Same-lang. 93±2% 84±2%
Cross-lang. 86±2% 74±3%
@@
Benchmark
User
Lang.
Doc
Lang.
Retrieval
Hit@20
End-to-End
Accuracy
Legal Arabic Arabic 87±4% 67±5%
Arabic English 51±5% 37±5%
@@
English English 88±4% 70±5%
Same-lang. 88±3% 69±4%
Cross-lang. 46±4% 29±3%
@@
Arabic English 54±4% 37±4%
English Arabic 64±4% 60±4%
English English 95±2% 85±3%
Same-lang. 92±2% 86±2%
Cross-lang. 59±3% 49±3%
@@
Table 2: Performance across language combinations. Results are presented for each embedder, benchmark and for each of the four possible user–document language combinations. In addition, we report same-language and cross-language scores, defined as the mean scores over combinations where the user and document languages match or differ, respectively. Each value is presented with its 95% confidence interval.
@@
The tables reveal that cross-lingual scenarios, where the user query and the supporting document are in different languages, consistently underperform compared to same-language settings in both retrieval and end-to-end performance. For the BGE-M3 embedder, a substantial decline in retrieval performance is observed only when the user language is English and the document language is Arabic, with drops of 33% in the Legal benchmark and 13% in Travel compared to the same-language configuration. A similar pattern appears in the final accuracy, with decreases of 37% and 14%, respectively. Notably, the reverse cross-lingual setting does not exhibit any statistically significant degradation for BGE-M3.
In contrast, the M-E5 embedding exhibits an even larger performance drop across both cross-lingual settings. Specifically, retrieval Hit@20 decreases by 42% on the Legal benchmark and by 33% on Travel, compared to their same-language counterparts. These retrieval declines also propagate to the end-to-end accuracy, resulting in drops of 40% for Legal and 37% for Travel.
@@ 4.3 The Source of the Cross-Lingual Failure
Notice that in our current setup, referred to from now on as the direct setting, we face two key challenges due to multilinguality. Firstly, “query-document language mismatch” requires the retriever to rank documents in one language in response to queries in another. Secondly, “document-document language mismatch” necessitates ranking documents across various languages without favoring high-resource languages or the user’s language.
To determine which of these challenges is primarily responsible for the observed failures, we conducted the following experiment. We modified the retriever from the direct setting to search only within the correct language. Specifically, for query corresponding to a (ground truth) document language X, the language-oracle retriever returns the top results exclusively in language X, completely excluding language Y. Hence, the language-oracle retriever has the "query-document language mismatch" challenge but completely avoids the "document-document language mismatch" challenge.
We stress that the oracle is used only for analysis purposes, since in practice we do not have access to the document language ahead of time. The first two bars in each subfigure of Figure 1 present the Hit@20 performance of the direct and language-oracle retrievers, broken down by query-document language combinations, as well as overall indicating the performance over the entire benchmark.
@@
We observe two clear phenomena. First, the language-oracle retriever achieves nearly identical performance across all query and document language pairs, suggesting there are essentially no failures related to the query-document language mismatch challenge. In contrast, the gap between the direct and language-oracle retrievers can be substantial in many cross-lingual cases. This indicates that the main source of failure lies in the document-document language mismatch challenge, namely the retriever’s ability to rank documents across languages.
The results suggests that while semantic similarity is well captured within a single language, the retrievers struggle in cross-lingual settings. For instance, BGE-M3 appears to favor English passages when the user query is in English, while M-E5 may exhibit a tendency to prefer passages in the same language as the user query.